OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line that is the perpendicular bisector of the segment whose endpoints are R(-1, 6) and S(5, 5)

OpenStudy (anonymous):

Do you know how to find the eq. of the line containing \(RS\)?

OpenStudy (anonymous):

eq?

OpenStudy (anonymous):

equation

OpenStudy (anonymous):

i think but i just don't know what to do when they say there is a perpendicular bisector

OpenStudy (anonymous):

The line you're looking for is \(perpendicular\) to \(RS\), and it also \(bisects\) it, which means the line you're looking for intersects \(RS\) at its midpoint.

OpenStudy (anonymous):

Perpendicular lines have slopes that are opposite reciprocals of each other, so the first thing you have to do is find the slope of the line containing \(RS\). Then you find its midpoint.

OpenStudy (anonymous):

so -6 is the slope

OpenStudy (anonymous):

Yep, so that means the slope of the perpendicular line is...

OpenStudy (anonymous):

well it was 1/6 so i did what you said to make it -6

OpenStudy (anonymous):

i got 1/6 for RS

OpenStudy (anonymous):

Sorry, you're right. I misread the points.

OpenStudy (anonymous):

all good :)

OpenStudy (anonymous):

so i end up with y=-6x what do i do from there? or is that even right?

OpenStudy (anonymous):

That's not the right equation, no. You have to use the point-slope formula: \[y-y_0=m(x-x_0)\] \((x_0,y_0)\) is either given point, and \(m\) is the slope. Plug in those values and solve for y.

OpenStudy (anonymous):

can you show me how to do that?

OpenStudy (anonymous):

Just a moment, I'm getting ahead of myself... Something doesn't look right to me.

OpenStudy (anonymous):

The slope of \(RS\) should be \(-\dfrac{1}{6}\), for one thing: \[\text{slope}=\frac{5-6}{5-(-1)}=\frac{-1}{6}\] which means the slope of the perp. bisector is actually \(6\).

OpenStudy (anonymous):

And my first instinct to use the point-slope formula was right, but I used the wrong info. The point I mention in the comment with the formula should be the midpoint of \(RS\). \[\text{midpoint}=\left(\frac{-1+5}{2},\frac{6+5}{2}\right)=\left(2,\frac{11}{2}\right)\] Now plug all this information into the point-slope formula. This is the equation of the line you're looking for: \[y-\frac{11}{2}=6(x-2)\] Do you know how to write the eq. of a line in standard form?

OpenStudy (anonymous):

no i'm not quite sure

OpenStudy (anonymous):

First thing you would do is distribute the 6, then move all the x's and y's to one side, and the constant to the other. Doing this, you have \[y-6x=-12+\frac{11}{2}\\ -6x+y=-\frac{13}{2}\] Standard form is \[Ax+By=C\]